1. One end of an iron poker is placed in a fire where the temperature is 502°C, and the other end is kept at a temperature of 26°C. The poker is 1.2m long and has a radius of 5.0×10^-3m. Ignoring the heat lost along the length of the poker, find the amount of heat conducted from one end of the poker to the other in 5.0s
1
Expert's answer
2018-08-20T10:12:09-0400
dQ/dt=kA(T_2-T_1 )/d
A=πr^2
The amount of heat conducted from one end of the poker to the other in 5.0s is
Q=dQ/dt ∆t=((79.5)π(0.005)^2 (502-26))/1.2 (5)=12 J.